Chloropropionic acid, CICH2CH2CooH is a weak monoprotic acid with Ka 7.94 x 10 5 M. Calculate the pH at the equivalence point in a titration 32.4 mL of 0.45 M chloropropionic acid with 0.4 M KoH.
Answer #1

at equivalnece point :

millimoles of acid = millimoles of base

32.4 x 0.45 = 0.4 x V

V = 36.45 mL

at equivalence point salt only reminas .

salt concentration = millimoles / total volume

= 14.58 / (32.4 + 36.45)

= 0.2118 M

pKa = - log Ka = 4.100

this is a salt of stong base and weak acid . so pH > 7

pH = 7 + 1/2 (pKa + log C)

= 7 + 1/2 (4.10 + log 0.2118)

= 8.71

pH = 8.71